# Learning Group Invariant Calabi–Yau Metrics by Fundamental Domain Projections

### v1

@@ 4.1.3 Scaling by roots of unity
and the projection map $h_{\rm rs}:X\to\mathcal{F}$ is given by [^19] We implement this projection in TensorFlow— by finding the angles with angles=tf.math.angle(complex_input)— and then applying the modulo operation with
@@
$h_{\rm rs}:\;[r_{0}e^{i\theta_{0}}:\ldots:r_{4}e^{i\theta_{4}}]\to[r_{0}\exp[i(\theta_{0}\mod 2\pi/5)],\ldots,r_{4}\exp[i(\theta_{4}\mod 2\pi/5)]]\,.$
@@
$h_{\rm rs}(\mathbf{z})=\rho(\pi(z))\mathbf{z}=\begin{pmatrix}\exp(2\pi i4/5)&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{pmatrix}\begin{pmatrix}r_{0}e^{2\pi i3/10}\\ r_{1}e^{2\pi i\varphi_{1}}\\ r_{2}e^{2\pi i\varphi_{2}}\\ r_{3}e^{2\pi i\varphi_{3}}\\ r_{4}e^{2\pi i\varphi_{4}}\end{pmatrix}=\begin{pmatrix}r_{0}e^{2\pi i1/10}\\ r_{1}e^{2\pi i\varphi_{1}}\\ r_{2}e^{2\pi i\varphi_{2}}\\ r_{3}e^{2\pi i\varphi_{3}}\\ r_{4}e^{2\pi i\varphi_{4}}\end{pmatrix}\,.$ (21)
@@ 4.2 Combining invariant layers
To achieve metric predictions that are invariant under a (semi)-direct product of groups, as is the case for the Fermat quintic, we would like to construct ML models that are invariant under several groups. This can be attained by adding the $G$-canonicalization layers in sequence to the network. However, we must then make sure that the canonicalizations satisfy a non-trivial condition: Let $G=G_{1}\ltimes G_{2}$ act on a space $X$ and let $h_{i}$ be a $G_{i}$-canonicalization for $i=1,2$. Then $h:=h_{2}\circ h_{1}$ defines a $G$-canonicalization if and only if $h_{1}$ descends to a well defined map on the orbits of $G$, namely $h_{1}:X/G_{2}\to X/G_{2}$ is well defined. Note that this condition on $h_{1}$ is weaker than $h_{1}$ being equivariant with respect to $G_{2}$.
To make this concrete, let’s return to the metric on the Fermat quintic $X$ defined by (3), which is invariant under action of the group $G=S_{5}\ltimes\mathbb{Z}_{5}^{5}$. For later reference, we note that the action of this group is not free. Let $h_{\rm perm}$ be the $S_{5}$-canonicalization defined ordering according norm, and let $h_{\rm rs}$ be a $\mathbb{Z}_{5}^{5}$-canonicalization as defined in (4.1.3). Then $h_{\rm rs}\circ h_{\rm perm}$ is a $G$-canonicalization, since for any $r\in\mathbb{Z}_{5}^{5}$ and $z\in X$ we have
@@
Performing a similar analysis, we find that $h_{\rm rs}\circ h_{\rm perm}\circ h_{\rm hom}$ is a [(×$[(\mathbb{C}^{\star}\times S_{5})\ltimes\mathbb{Z}_{5}^{5}]$-canonicalization, where $h_{\rm hom}$ is given by (4.1.1), and $\text{Spectral}\circ h_{\rm perm}$ is invariant under $S_{5}\times\mathbb{C}^{\star}$.
@@ 5.1 Fermat quintic
Figure 3: Performance of different ML models on the Fermat quintic. From left to right: Monge–Ampère, Ricci, and transition loss evolution on validation data; volume computed with trained metric on all data (ground truth is 5). The upper (lower) row shows dense (spectral) networks with and without G-canonicalization layers. Solid line is average over 5 experiments, lighter color bands are the standard deviation.
We have performed a number of experiments on this geometry, and report the results from training the models on 90 000 points, over 50 epochs in figure 3 (we have checked that the conclusions are reproduced in 300-epochs experiments). Since training and validation plots look very similar, we only report the latter. [^21] In more detail: the MA and transition losses on training data look very similar to the shown validation plots, showing that there is no problem with overfitting. The Kclass loss evolution is consistent with the Volume plot on validation data. Kähler and Ricci loss is disabled during training. Thus, figure 3 shows the evolution, over the training epochs, of Monge-Ampère, Ricci and transition loss on a validation set consisting of 10 000 points (the Kähler loss evaluates to $\mathcal{O}(10^{-16})$, and is omitted from the plot). We also track the evolution of the volume, as computed by the integral of the volume measure of the predicted metric (cf. appendix B). Since the computation of the volume requires an integral, it is performed over the full data set, i.e. 100 $0$ points.
@@
When examining the Ricci loss, we notice a number of interesting features. First, the spectral layer has a larger positive impact on performance, both at initialization and during training. Second, while the canonicalization layers may have a positive impact on the Ricci loss at initialization of standard $\phi$-models, the effect becomes negative after 10 epochs of training. While this discrepancy decreases with training, the standard $\phi$-model has a slightly lower Ricci loss on validation than its canonicalized counterparts. For spectral models, adding canonicalized layers has an even larger negative effect on the Ricci loss. We interpret this as an effect of a complicated loss landscape where optimizing with respect to the MA loss does not necessarily improve the Ricci loss (despite the fact that both of these losses are zero for the true Ricci-flat metric). Similar effects have been observed when learning CY metrics on more complicated CICY manifolds [40], and may here be related to the fact that the symmetry group has fixed points, that may affect the computation of Ricci loss. As suggested in ref. [40], by making a different hyperparameter choice, so that both MA and Ricci losses are used to guide the training of the ML model, performance may be improved; we leave such studies for the future.
@@ 6 Ricci-flat metrics on non simply connected CY manifolds
However, it is somewhat non-trivial to determine the basis of sections on $Q$, i.e. identify the linearly indepentdent $G$-invariant sections on $X$ that restrict to $Q$, already for moderate values of $k$. There are two subtleties: first, polynomials of the form $q_{i}p$, where $p=0$ defines the CY hypersurface, evaluate to zero on the CY, and must be removed from the invariant basis. Second, the generators $g_{1},g_{2}$ only commute up to homogeneous rescaling of the coordinates, which makes the construction of invariant polynomials non-trivial. Braun and collaborators rely on algorithms of invariant theory, implemented using SINGULAR [60], to identify the basis, and explore these methods up to $k=25$. After running the Donaldson algorithm on a a sample with $10^{6}$ points, they find that the $\sigma$ measure evaluates to $0.015$ for $k=25$ on an unseen test set of 20 000 points.242424 When comparing the performance against number of basis elements, they find that the $G$-invariant metric approximation has smaller $\sigma$ error than a metric computed using the unrestricted spectral basis.
@@
We implement the canonicalization layers as follows. Let $G_{i}\subset G$ be the subgroup generated by the generator $g_{i}$ for $i=1,2$. Then, by projecting on the fundamental domains of $G_{1}$ and $G_{2}$ consecutively, we construct an invariant function on $X$ w.r.t $G$ or, in other words, a function on the smooth quotient $Q=X/G$. A technical complication, compared to the Fermat quintic canonicalization layers of Sec. 5.1 is that we need to address the non-commutativity of $G_{1}$ and $G_{2}$. This can be resolved by inserting the homogeneous canonicalization layer before the other $G$-canonicalization layers; $g_{1}$ and $g_{2}$ commute up to homogeneous rescalings.
@@
To project on the fundamental domains $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, we implement two canonicalization layers that act onn the homogeneous coordinates $[z_{0}:z_{1}:z_{2}:z_{3}:z_{4}]$ of a point on $X$. The first layer returns the coordinate tuple shifted to the left such that the highest norm entry takes the first spot,
@@ 7 Conclusion
Our work is inspired by recent developments in the ML literature, that advocate to achieve model invariance by a pre-processing step, where the input data is projected into a fundamental domain of the relevant group [35, 36]. We have used an approach where non-trainable layers performs these projections; collectively, we refer to these layers as canonicalization layers. This is a modular implementation, and the layers can easily be exchanged to match different setting. While we have limited our studies to the cymetric package [21, 22], we expect similar canonicalization layers can be added with ease in other ML packages for CY geometry.
We have performed a number of experiments on different CICY manifolds, comparing the performance of the standard $\phi$-model of the cymetric package with and without canonicalization layers, but otherwise identical hyperparameter settings. We have also compared with the spectral network, originally proposed by Berglund and collaborators [24], which may also be combined with canonicalization layers. We find that adding canonicalization layers increases performance on the losses that are used to train the networks. In particular, the increase in performance on the transition loss is several orders of magnitudes for all architectures. The volume, controlled by the K-class loss, converges faster if canonicalization layers are added to the standard $\phi$-model. However, the volume seems to be captured best by the spectral networks without canonicalization. Being a global property, this could be explained by discontinuities, related to the boundaries of fundamental domains, that the canonicalization layers may introduce. These are not present in the spectral layer.
@@ Acknowledgement
We thank Lara Anderson, Georgios Dimitroglou Rizell, Cristofero Fraser-Taliente, James Halverson, Thomas Hardy, Carlos Rodriguez, Fabian Ruehle, Robin Schneider, and Ayca Özcelikkale for valuable discussions and suggestions. We also thank Carlos Rodriguez and Robin Schneider for comments on a draft of this paper; Fabian Ruehle for his continuous work in maintaining the cymetric package; Muyang Liu and Martin Svärdsjö for collaborations on related topics; and, finally, Daniel Persson and Jan Gerken for organizing the Mathy-AI Journal Club, as well as its attendants, for providing a lot of inspiration for this work.

### v2

@@ 4.1.3 Scaling by roots of unity
and the projection map $h_{\rm RootSc}:X\to\mathcal{F}$ is given by [^20] We implement this projection in TensorFlow— by finding the angles with angles=tf.math.angle(complex_input)— and then applying the modulo operation with
@@
$h_{\rm RootSc}:\;[r_{0}e^{i\theta_{0}}:\ldots:r_{4}e^{i\theta_{4}}]\to[r_{0}\exp[i(\theta_{0}\mod 2\pi/5)],\ldots,r_{4}\exp[i(\theta_{4}\mod 2\pi/5)]]\,.$
@@
$h_{\rm RootSc}(\mathbf{z})=\rho(\pi(z))\mathbf{z}=\begin{pmatrix}\exp(2\pi i4/5)&0&0&0&0\\ 0&1&0&0&0\\ 0&0&1&0&0\\ 0&0&0&1&0\\ 0&0&0&0&1\end{pmatrix}\begin{pmatrix}r_{0}e^{2\pi i3/10}\\ r_{1}e^{2\pi i\varphi_{1}}\\ r_{2}e^{2\pi i\varphi_{2}}\\ r_{3}e^{2\pi i\varphi_{3}}\\ r_{4}e^{2\pi i\varphi_{4}}\end{pmatrix}=\begin{pmatrix}r_{0}e^{2\pi i1/10}\\ r_{1}e^{2\pi i\varphi_{1}}\\ r_{2}e^{2\pi i\varphi_{2}}\\ r_{3}e^{2\pi i\varphi_{3}}\\ r_{4}e^{2\pi i\varphi_{4}}\end{pmatrix}\,.$ (21)
@@ 4.2 Combining invariant layers
To achieve metric predictions that are invariant under a (semi)-direct product of groups, as is the case for the Fermat quintic, we would like to construct ML models that are invariant under several groups. This can be attained by adding the $G$-canonicalization layers in sequence to the network. However, we must then make sure that the canonicalizations satisfy a non-trivial condition: Let $G=G_{1}\ltimes G_{2}$ act on a space $X$ and let $h_{i}$ be a $G_{i}$-canonicalization for $i=1,2$. Then $h:=h_{2}\circ h_{1}$ defines a $G$-canonicalization if and only if $h_{1}$ descends to a well defined map on the orbits of $G_{2}$, namely $h_{1}:X/G_{2}\to X/G_{2}$ is well defined. Note that this condition on $h_{1}$ is weaker than $h_{1}$ being equivariant with respect to $G_{2}$.
To make this concrete, let’s return to the metric on the Fermat quintic $X$ defined by (3), which is invariant under action of the group $G=S_{5}\ltimes\mathbb{Z}_{5}^{5}$. For later reference, we note that the action of this group is not free. Let $h_{\rm perm}$ be the $S_{5}$-canonicalization defined by (17), and let $h_{\rm RootSc}$ be a $\mathbb{Z}_{5}^{5}$-canonicalization as defined in (4.1.3). Then $h_{\rm RootSc}\circ h_{\rm perm}$ is a $G$-canonicalization, since for any $r\in\mathbb{Z}_{5}^{5}$ and $z\in X$ we have
@@
Performing a similar analysis, we find that $h_{\rm RootSc}\circ h_{\rm perm}\circ h_{\rm hom}$ is a [(×$[(\mathbb{C}^{\star}\times S_{5})\ltimes\mathbb{Z}_{5}^{5}]$-canonicalization, where $h_{\rm hom}$ is given by (4.1.1), and $\text{Spectral}\circ h_{\rm perm}$ is invariant under $S_{5}\times\mathbb{C}^{\star}$.
@@ 5.1 Fermat quintic
Figure 3: Performance of different ML models on the Fermat quintic. From left to right: Monge–Ampère, Ricci, and transition loss evolution on validation data; volume computed with trained metric on all data (ground truth is 5). The upper (lower) row shows dense (spectral) networks with and without G-canonicalization layers. Solid line is average over 5 experiments, lighter color bands show the standard deviation.
We have performed a number of experiments on this geometry, and report the results from training the models on 90 000 points, over 50 epochs in figure 3 (we have checked that the conclusions are reproduced in 300-epochs experiments). Since training and validation plots look very similar, we only report the latter. [^22] In more detail: the MA and transition losses on training data look very similar to the shown validation plots, showing that there is no problem with overfitting. The Kclass loss evolution is consistent with the Volume plot on validation data. Kähler and Ricci loss are disabled during training. Thus, figure 3 shows the evolution, over the training epochs, of Monge-Ampère, Ricci and transition loss on a validation set consisting of 10 000 points (the Kähler loss evaluates to $\mathcal{O}(10^{-16})$, and is omitted from the plot). We also track the evolution of the volume, as computed by the integral of the volume measure of the predicted metric (cf. appendix B). Since the computation of the volume requires an integral, it is performed over the full data set, i.e. 100 $0$ points.
@@
When examining the Ricci loss, we notice a number of interesting features. First, the spectral layer has a larger positive impact on performance, both at initialization and during training. Second, while the canonicalization layers may have a positive impact on the Ricci loss at initialization of standard $\phi$-models, the effect becomes negative after 10 epochs of training. While this discrepancy decreases with training, the standard $\phi$-model has a slightly lower Ricci loss on validation than its canonicalized counterparts. For spectral models, adding canonicalized layers has a negative effect on the Ricci loss. We interpret this as an effect of a complicated loss landscape where optimizing with respect to the MA loss does not necessarily improve the Ricci loss (despite the fact that both of these losses are zero for the true Ricci-flat metric). Similar effects have been observed when learning CY metrics on more complicated CICY manifolds [40], and may here be related to the fact that the symmetry group has fixed points, that may affect the computation of Ricci loss. As suggested in ref. [40], by making a different hyperparameter choice, so that both MA and Ricci losses are used to guide the training of the ML model, performance may be improved; we leave such studies for the future.
@@ 6 Ricci-flat metrics on non simply connected CY manifolds
However, it is somewhat non-trivial to determine the basis of sections on $Q$, i.e. identify the linearly independent $G$-invariant sections on $X$ that restrict to $Q$, already for moderate values of $k$. There are two subtleties: first, polynomials of the form $q_{i}p$, where $p=0$ defines the CY hypersurface, evaluate to zero on the CY, and must be removed from the invariant basis. Second, the generators $g_{1},g_{2}$ only commute up to homogeneous rescaling of the coordinates, which makes the construction of invariant polynomials non-trivial. Braun and collaborators rely on algorithms of invariant theory, implemented using SINGULAR [60], to identify the basis, and explore these methods up to $k=25$. After running the Donaldson algorithm on a a sample with $10^{6}$ points, they find that the $\sigma$ measure evaluates to $0.015$ for $k=25$ on an unseen test set of 20 000 points.252525 When comparing the performance against number of basis elements, they find that the $G$-invariant metric approximation has smaller $\sigma$ error than a metric computed using the unrestricted spectral basis.
@@
We implement the canonicalization layers as follows. Let $G_{i}\subset G$ be the subgroup generated by the generator $g_{i}$ for $i=1,2$. Then, by projecting on the fundamental domains of $G_{1}$ and $G_{2}$ consecutively, we construct an invariant function on $X$ w.r.t $G$ or, in other words, a function on the smooth quotient $Q=X/G$. A technical complication, compared to the Fermat quintic canonicalization layers of Sec. 5.1, is that we need to address the non-commutativity of $G_{1}$ and $G_{2}$. This can be resolved by inserting the homogeneous canonicalization layer before the other $G$-canonicalization layers; $g_{1}$ and $g_{2}$ commute up to homogeneous rescalings.
@@
To project on the fundamental domains $\mathcal{F}_{1}$ and $\mathcal{F}_{2}$, we implement two canonicalization layers that act on the homogeneous coordinates $[z_{0}:z_{1}:z_{2}:z_{3}:z_{4}]$ of a point on $X$. The first layer returns the coordinate tuple shifted to the left such that the highest norm entry takes the first spot,
@@ 7 Conclusion
Our work is inspired by recent developments in the ML literature, that advocate to achieve model invariance by a pre-processing step, where the input data is projected into a fundamental domain of the relevant group [35, 36]. We have used an approach where non-trainable layers perform these projections; collectively, we refer to these layers as canonicalization layers. This is a modular implementation, and the layers can easily be exchanged to match different settings. While we have limited our studies to the cymetric package [21, 22], we expect similar canonicalization layers can be added with ease in other ML packages for CY geometry.
We have performed experiments on different CICY manifolds, comparing the performance of the standard $\phi$-model of the cymetric package with and without canonicalization layers, but otherwise identical hyperparameter settings. We have also compared with the spectral network, originally proposed by Berglund and collaborators [24], which may also be combined with canonicalization layers. We find that adding canonicalization layers increases performance on the losses that are used to train the networks. In particular, the increase in performance on the transition loss is several orders of magnitudes for all architectures. The volume, controlled by the K-class loss, converges faster if canonicalization layers are added to the standard $\phi$-model. However, the volume seems to be captured best by the spectral networks without canonicalization. Being a global property, this could be explained by discontinuities, related to the boundaries of fundamental domains, that the canonicalization layers may introduce. These are not present in the spectral layer.
@@ Acknowledgement
We thank Lara Anderson, Georgios Dimitroglou Rizell, Cristofero Fraser-Taliente, James Halverson, Thomas Harvey, Daniel Platt, Carlos Rodriguez, Fabian Ruehle, Robin Schneider, and Ayca Özcelikkale for valuable discussions and suggestions. We also thank Carlos Rodriguez and Robin Schneider for comments on a draft of this paper; Fabian Ruehle for his continuous work in maintaining the cymetric package; Muyang Liu and Martin Svärdsjö for collaborations on related topics; and, finally, Daniel Persson and Jan Gerken for organizing the Mathy-AI Journal Club, as well as its attendants, for providing a lot of inspiration for this work.